OpenStudy (anonymous):

a charge q is placed at the center of the line joining two equal charges Q. shoW that the system of three charge will be in equilibrium if q= -Q/4

OpenStudy (anonymous):

 assume that q is positive Get the total force on Q and make it equal to zero as it is in equilibrium

OpenStudy (anonymous):

Let us treat left Q charge as the test charge and other two charges q and Q as source charges. For equilibrium force acting on the test charge should be zero. \[Force = k\frac{q_1 q_2}{r^2}\] Here, \[Force \ between \ q \ and \ Q = k\frac{qQ}{(r/2)^2}\] Here, \[Force \ between \ Q \ and \ Q = k\frac{Q^2}{(r)^2}\] Total force = 0 So \[k\frac{qQ}{(r/2)^2} +k\frac{Q^2}{(r)^2} = 0 \] Solving this equation gives q = -Q/4